OpenStudy (perii224):

PLEASE HELP WILL MEDAL The sum of two numbers is 66. The larger number is 2 more than 3 times the smaller. What are the two numbers? 22 and 44 20 and 46 18 and 48 16 and 50 Suppose the amount of money in a bank account with an initial balance of $25,000 earns 1.5 percent interest per year. If the exponential function, y=AxB^x is used to model this situation, which value would represent B in the function? 25,000 1.15 1.5 1.015

OpenStudy (perii224):

@Evoker

OpenStudy (evoker):

Lets call the numbers x and y with x being the larger, then your first statement says that x+y=66 and the second says that x=2+3y. Substituting into the first equation the result from the second equation gives 2+3y+y=66 or 2+4y=66 which yields y=16, then using any of the previous equations we get that x=50 so 16 and 50 are your two numbers.

OpenStudy (zzr0ck3r):

x+y=66 3x+2=y

OpenStudy (evoker):

for the second question B would be the interest rate added to 1 or 1.015 since 1.5 percent in decimal is .015

OpenStudy (robtobey2):

{a + b = 66, 3 a + 2 = b} // Solve {a = 16, b = 50}

OpenStudy (perii224):

Thank you all so much! :]